# Solving the Correlation Cluster LP in Sublinear Time

### v2

@@ 7.1 Compute $d_{\mathrm{cross}}(v)$
$\Pr\left[\left|\frac{s}{m_{K}}\leavevmode\nobreak\ X-\frac{s}{m_{K}}\leavevmode\nobreak\ |E^{+}(K,V\setminus K)|\right|\geq\frac{\beta}{4}\cdot\frac{s}{m_{K}}\leavevmode\nobreak\ |E^{+}(K,V\setminus K)|\right]$ $\leq 2\exp\left(-\frac{\beta^{2}}{16}\cdot\frac{s}{m_{K}}\leavevmode\nobreak\ |E^{+}(K,V\setminus K)|\right)$
@@
Since $d_{\mathrm{cross}}(K)=|E^{+}(K,V\setminus K)|+2|E^{-}(K)|\leq\beta|K|$, we know that $|E^{+}(K,V\setminus K)|\leq\beta|K|$ and $|E^{-}(K)|\leq\frac{1}{2}\beta|K|$. Thus, with high probability, $X\leq 3\beta|K|$ and $Y\leq\frac{3}{2}\beta|K|$. Hence, $\hat{d}_{\mathrm{cross}}(K)=X+2Y\leq 6\beta|K|$Diamond
@@ 7.4 Finding one small ratio cluster in sublinear time
8:  $\hat{D}(r)\leftarrow\hat{N}_{\mathrm{cand}}(r)\setminus{K(r)}$
9:  $\mathcal{T}\leftarrow\emptyset$
@@
12:        Add $T\leftarrow$ GenerateCluster($r,D(r),S^{1},\ldots,S^{\eta},\tilde{t}_{1},\ldots,\tilde{t}_{\eta}$) to $\mathcal{T}$.
@@
Algorithm 8 GenerateCluster($r,\hat{D}(r),S^{1},\ldots,S^{\eta},\tilde{t}_{1},\ldots,\tilde{t}_{\eta}$)
1:  $T\leftarrow K(r),\hat{T}_{1}\leftarrow K(r)$
@@ 7.5 Determine the correct guess $R$ of the optimal cost
Our final challenge is to determine the best solution $z$ among the solutions we computed for each guess $R$. Since we iterate over different granularities of $R$ in the range $[1,n^{2}]$, we know that one of these values satisfies $R\in[{\mathrm{cover}}(\mathrm{OPT}),(1+\gamma){\mathrm{cover}}(\mathrm{OPT})]$. Thus, we know that for this guess $R$, we will compute a solution $z$ with small cost. To determine the this solution (or any other with a smaller cost), we will use Lemma 48 to estimate the cost. After applying Algorithm 1, we obtain a solution $\{z_{S}\}$, where each nonzero $z_{S}$ is at least some constant. We can estimate $\mathrm{cost}(S)$ for each nonzero $z_{S}$, and since each node appears in only a constant number of sets $S$, the total running time for estimating all $\mathrm{cost}(S)$ values is $\tilde{O}(n)$.
@@ 8 MPC Implementation
4:     If $\frac{{\mathrm{cover}}(K(v))}{p(K(v))}\leq(1+6\gamma)\leavevmode\nobreak\ R$, add $K(v)$ to $\mathcal{F}$ and set $\hat{q}_{w}=0$ for all $w\in K(v)$.
@@ 8.1 Approximate Ratio of Algorithm 9
Given vertex weights $p_{v}>0$, Algorithm 9 finds a family =$\mathcal{F}=\{S_{1},S_{2},...,S_{l}\mid S_{i}\subseteq V\}$ such that,
@@
    for any distinct $S,T\in\mathcal{F}$, $S\cap T=\emptyset$,
@@ 8.2 Number of Iterations of Algorithm 9
and $C\cap U\not=\emptyset$ with probability at least
@@
Notice that the algorithm will set some $\hat{q}_{v}$ to 00 without adding $u$ to $\mathcal{F}$, let
@@
For the second inequality, we use that $\gamma\leavevmode\nobreak\ {\mathrm{cover}}(\mathrm{OPT})\geq\gamma\leavevmode\nobreak\ \mathrm{cost}(\mathrm{OPT})\geq\gamma^{2}|E_{\text{adm}}|$ because of the preclustering (Theorem 7). On the other hand, if $\hat{q}(\mathcal{C}_{\leq(1+4\gamma)R})<\gamma$, then consider $C\in\mathrm{OPT}$ whose ratio is at least $(1+4\gamma)R$, we have
@@
$\mathcal{C}_{\leq(1+4\gamma)R}=\{C\in\mathcal{C}^{*}_{\leq(1+4\gamma)R}\mid C\cap\widetilde{U}\not=\emptyset\}$
@@
when we set up $\hat{p}$, we first copy $\hat{q}$ to $\hat{p}$ value. Then, we might set some $\hat{p}$ value to 00 if the nodes have two chosen neighbor. So, to show that $\hat{p}(C^{*})\geq(1-\gamma)\hat{q}(C^{*})$, we need to show that the removed nodes contributes to at most $\gamma$ fractional weight to $C^{*}$.
@@
Now for each $C^{*}\in\mathcal{C}^{*}_{\leq(1+4\gamma)R}$, based on Claim 52, we know that with probability at least $\Omega(\varepsilon^{5}\gamma^{\prime})$, $C^{*}\cap\widetilde{U}\not=\emptyset$ and $C^{*}\in\mathcal{C}_{\leq(1+4\gamma)R}$, so we have
@@ 8.3 Wrap-Up: Proof of MPC Algorithm for Theorem 1
Finally, we need to apply Lemma 11 to compute the solution for cluster LP. Implementing Lemma 11 in the MPC model is somewhat tricky, but the key observation is that each $z_{S}\geq\frac{1}{T_{\mathrm{MW}}}$. Therefore, for each node $u$, at most $O(T_{\mathrm{MW}})$ clusters with positive $z_{S}$ values will cover $u$. For each node, we collect all clusters with positive $z_{S}$ values. If the final sum $\sum_{S\ni u}z_{S}>1$, we sort $z_{S}$ for $u$ in any order and mark those $z_{S}$ as ”greater than 1 for $u$” for later clusters. Each $S$ then collects these marked clusters and removes them from itself. In this process, for each non-singleton atom, we perform this operation only once for a representative, allowing $S$ to remove the non-singleton atoms. Since at most $O(T_{\mathrm{MW}})$ clusters cover $u$, this process requires at most $O(T_{\mathrm{MW}}\cdot n)$ total space.
@@ 9 Rounding Algorithms
[11] There exists an algorithm that, given a feasible solution for the cluster LP, produces a clustering whose objective value is at most $1.437$ times that of the cluster LP solution.
To be more precise, the algorithm from Theorem 56 consists of two different rounding algorithms, one is executed with probability $p=\frac{1.437}{2}$ the other with probability $1-p$. The two algorithms in question are the cluster-based rounding (Algorithm 10) and the pivot-based rounding (Algorithm 11). We need to show that we can implement both algorithms in nearly linear time.
@@
The pivot-based algorithm (Algorithm 11) follows a different approach. Instead of selecting a set directly, at each round, we randomly choose a node $u$ that is still in the graph, referred to as the pivot. We then form a cluster using this pivot as follows. For each small $+$edge $(u,v)$, we include $v$ in the cluster. For the remaining $+$edges, we apply correlated rounding, where we randomly choose a set $S$ containing $u$ and include all remaining $+$edges if they are in $S$. For --edges, we use independent rounding, where the probability of adding a node to the cluster is determined by the distance metric:
@@
Algorithm 11 Pivot-Based Rounding with Threshold $1/3$
1:  $\mathcal{C}\leftarrow\emptyset,V^{\prime}\leftarrow V$
@@
3:     randomly choose a pivot $u\in V^{\prime}$
4:     $C\leftarrow\{v\in V^{\prime}\cap N^{+}(u):x_{uv}\leq\frac{1}{3}\}$
5:     for every $v\in V^{\prime}\cap N^{-}(u)$ do
6:        independently add $v$ to $C$ with probability $1-x_{uv}$
@@
8:     randomly choose a set $S\ni u$, with probabilities $z_{S}$We have $\sum_{S\ni u}z_{S}=1$}
9:     $C\leftarrow C\cup(S\cap V^{\prime}\cap N^{+}(u))$, $\mathcal{C}\leftarrow\mathcal{C}\cup\{C\}$, $V^{\prime}\leftarrow V^{\prime}\setminus C$
10:  end while
11:  return $\mathcal{C}$
@@ 9.1 Nearly Linear Time Rounding Algorithm
Now we discuss how to implement Algorithm 11 in nearly linear time. Note that each vertex is contained in at most $\frac{1}{\Delta}$ clusters $S\in\mathrm{supp}(z)$. Thus a fixed pivot $u$ belongs to at most $1/\Delta$ sets $S$ such that $z_{S}>0$. We therefore only need to consider $1/\Delta$ sets to choose one of these sets according to the respective probability distribution. Visiting the $+$edges incident to each pivot requires at most $O(m)$ time over the course of the algorithm.
For the --edges, the main challenge is listing all --edges from $V\cap N^{-}(u)$ and then performing independent rounding. Our key observation is that $x_{uv}<1$ if and only if there exists some $S$ such that $S$ contains both $u$ and $v$ and $z_{S}>0$. Thus, we only need to iterate over all vertices in $\cup_{S\ni u,z_{S}>0}\leavevmode\nobreak\ S$.
Since each $z_{S}\geq\Delta$ for $S\in\mathrm{supp}(z)$, each node $v$ with $x_{uv}<1$ contributes at least $1/\Delta$ to the LP value. Therefore, there are at most $O(m/\Delta)$-edges across all $S$. Since each --edge is visited at most once, listing all --edges takes at most $O(m/\Delta)$ time.
@@ 9.2 Rounding in Sublinear Model
On the other hand, for each vertex $v\in U_{r}$, the probability that it is included in the cluster of $r$ is at least $1-x_{rv}\geq\Delta$. Thus, in expectation, we remove at least $\Delta|U_{r}|$ vertices from the current graph.

### v5

@@ 7.1 Compute $d_{\mathrm{cross}}(v)$
$\Pr\left[\left|\frac{s}{m_{K}}\penalty 10000\ X-\frac{s}{m_{K}}\penalty 10000\ |E^{+}(K,V\setminus K)|\right|\geq\frac{\beta}{4}\cdot\frac{s}{m_{K}}\penalty 10000\ |E^{+}(K,V\setminus K)|\right]$ $\leq 2\exp\left(-\frac{\beta^{2}}{16}\cdot\frac{s}{m_{K}}\penalty 10000\ |E^{+}(K,V\setminus K)|\right)$
@@
Since $d_{\mathrm{cross}}(K)=|E^{+}(K,V\setminus K)|+2|E^{-}(K)|\leq\beta|K|$, we know that $|E^{+}(K,V\setminus K)|\leq\beta|K|$ and $|E^{-}(K)|\leq\frac{1}{2}\beta|K|$. Thus, with high probability, $X\leq 3\beta|K|$ and $Y\leq\frac{3}{2}\beta|K|$. Hence, $\hat{d}_{\mathrm{cross}}(K)=X+2Y\leq 6\beta|K|$ $\Diamond$
@@ 7.4 Finding one small ratio cluster in sublinear time
[…]
@@
[…]
@@
[…]
@@ 7.5 Determine the correct guess $R$ of the optimal cost
Our final challenge is to determine the best solution $z$ among the solutions we computed for each guess $R$. Since we iterate over different granularities of $R$ in the range $1,n^{2}]$, we know that one of these values satisfies $R\in[{\mathrm{cover}}(\mathrm{OPT}),(1+\gamma){\mathrm{cover}}(\mathrm{OPT})]$. Thus, we know that for this guess $R$, we will compute a solution $z$ with small cost. To determine the this solution (or any other with a smaller cost), we will use Lemma 48 to estimate the cost. After applying Algorithm 1, we obtain a solution $\{z_{S}\}$, where each nonzero $z_{S}$ is at least some constant. We can estimate $\mathrm{cost}(S)$ for each nonzero $z_{S}$, and since each node appears in only a constant number of sets $S$, the total running time for estimating all $\mathrm{cost}(S)$ values is $\tilde{O}(n)$.
@@ 8 MPC Implementation
4:  If $\frac{{\mathrm{cover}}(K(v))}{p(K(v))}\leq(1+6\gamma)\penalty 10000\ R$, add $K(v)$ to $\mathcal{F}$ and set $\hat{q}_{w}=0$ for all $w\in K(v)$.
@@ 8.1 Approximate Ratio of Algorithm 9
Given vertex weights $p_{v}>0$, Algorithm 9 finds a family $\mathcal{F}=\{S_{1},S_{2},...,S_{l}\mid S_{i}\subseteq V\}$ such that,
@@
    for any distinct $\in\mathcal{F}, ST=S\cap T$$\emptyset$,
@@ 8.2 Number of Iterations of Algorithm 9
and $C\cap U\not$$\emptyset$ with probability at least
@@
Notice that the algorithm will set some $\hat{q}_{v}$ to 0 without adding $u$ to $\mathcal{F}$, let
@@
For the second inequality, we use that $\gamma\penalty 10000\ {\mathrm{cover}}(\mathrm{OPT})\geq\gamma\penalty 10000\ \mathrm{cost}(\mathrm{OPT})\geq\gamma^{2}|E_{\text{adm}}|$ because of the preclustering (Theorem 7). On the other hand, if $\hat{q}(\mathcal{C}_{\leq(1+4\gamma)R})<\gamma$, then consider $C\in\mathrm{OPT}$ whose ratio is at least $(1+4\gamma)R$, we have
@@
$\mathcal{C}_{\leq(1+4\gamma)R}=\{C\in\mathcal{C}^{*}_{\leq(1+4\gamma)R}\mid C\cap\widetilde{U}\not$$\emptyset\}$
@@
when we set up $\hat{p}$, we first copy $\hat{q}$ to $\hat{p}$ value. Then, we might set some $\hat{p}$ value to 0 if the nodes have two chosen neighbor. So, to show that $\hat{p}(C^{*})\geq(1-\gamma)\hat{q}(C^{*})$, we need to show that the removed nodes contributes to at most $\gamma$ fractional weight to $C^{*}$.
@@
Now for each $C^{*}\in\mathcal{C}^{*}_{\leq(1+4\gamma)R}$, based on Claim 52, we know that with probability at least $\Omega(\varepsilon^{5}\gamma^{\prime})$, $C^{*}\cap\widetilde{U}\not$$\emptyset$ and $C^{*}\in\mathcal{C}_{\leq(1+4\gamma)R}$, so we have
@@ 8.3 Wrap-Up: Proof of MPC Algorithm for Theorem 1
Finally, we need to apply Lemma 11 to compute the solution for cluster LP. Implementing Lemma 11 in the MPC model is somewhat tricky, but the key observation is that each $z_{S}\geq\frac{1}{T_{\mathrm{MW}}}$. Therefore, for each node $u$, at most $O(T_{\mathrm{MW}})$ clusters with positive $z_{S}$ values will cover $u$. For each node, we collect all clusters with positive $z_{S}$ values. If the final sum $\sum_{S\ni u}z_{S}>1$, we sort $z_{S}$ for $u$ in any order and mark those $z_{S}$ as ”greater than 1 for uu” for later clusters. Each $S$ then collects these marked clusters and removes them from itself. In this process, for each non-singleton atom, we perform this operation only once for a representative, allowing $S$ to remove the non-singleton atoms. Since at most $O(T_{\mathrm{MW}})$ clusters cover $u$, this process requires at most $O(T_{\mathrm{MW}}\cdot n)$ total space.
@@ 9 Rounding Algorithms
[11] There exists an algorithm that, given a feasible solution for the cluster LP, produces a clustering whose objective value is at most $1.485$ times that of the cluster LP solution.
To be more precise, the algorithm from Theorem 56 consists of two different rounding algorithms, one is executed with probability $p=\frac{1.485}{2}$ the other with probability 1p1-p. The two algorithms in question are the cluster-based rounding (Algorithm 10) and the pivot-based rounding (Algorithm 11). We need to show that we can implement both algorithms in nearly linear time.
@@
The pivot-based algorithm (Algorithm 11) follows a different approach. Instead of selecting a set directly, at each round, we randomly choose a node $u$ that is still in the graph, referred to as the pivot. We then form a cluster using this pivot as follows. For each small ++edge $(u,v)$, we include $v$ in the cluster. For the remaining ++edges, we apply correlated rounding, where we randomly choose a set $S$ containing $u$ and include all remaining ++edges if they are in $S$. For -edges, we use independent rounding, where the probability of adding a node to the cluster is determined by the distance metric:
@@
Algorithm 11 Pivot-Based Rounding (with correlated mid-range for ++ edges)
[…]
@@
3:  choose a pivot $u\in V^{\prime}$ uniformly at random
4:  $C\leftarrow\{u\}$
5:  for each $v\in V^{\prime}\cap N^{-}(u)$ do
6:   independently add $v$ to $C$ with probability $1-x_{uv}^{2}$
@@
8:  $C_{\ell}\leftarrow\{v\in V^{\prime}\cap N^{+}(u):x_{uv}\leq 0.40\}$
9:  $C_{m}\leftarrow\{v\in V^{\prime}\cap N^{+}(u):0.40<x_{uv}\leq 0.57\}$
10:  $C_{h}\leftarrow\{v\in V^{\prime}\cap N^{+}(u):x_{uv}>0.57\}$
11:  for each $v\in C_{h}$ do
12:   independently add $v$ to $C$ with probability $1-x_{uv}$
13:  end for
14:  sample a random set $S\ni u$ according to the set distribution $\{z_{S}\}_{S\ni u}$
15:  $C\leftarrow C\cup C_{\ell}\cup\bigl(C_{m}\cap S\bigr)$
16:  $\mathcal{C}\leftarrow\mathcal{C}\cup\{C\}$, $V^{\prime}\leftarrow V^{\prime}\setminus C$
17:end while
18:return $\mathcal{C}$
@@ 9.1 Nearly Linear Time Rounding Algorithm
Now we discuss how to implement Algorithm 11 in nearly linear time. Note that each vertex is contained in at most $\frac{1}{\Delta}$ clusters $S\in\mathrm{supp}(z)$. Thus a fixed pivot $u$ belongs to at most $1/\Delta$ sets $S$ such that $z_{S}>0$. We therefore only need to consider $1/\Delta$ sets to choose one of these sets according to the respective probability distribution. Visiting the ++edges incident to each pivot requires at most $O(m)$ time over the course of the algorithm.
For the -edges, the main challenge is listing all -edges from $V\cap N^{-}(u)$ and then performing independent rounding. Our key observation is that $x_{uv}<1$ if and only if there exists some $S$ such that $S$ contains both $u$ and $v$ and $z_{S}>0$. Thus, we only need to iterate over all vertices in $\cup_{S\ni u,z_{S}>0}\penalty 10000\ S$.
Since each $z_{S}\geq\Delta$ for $S\in\mathrm{supp}(z)$, each node $v$ with $x_{uv}<1$ contributes at least $1/\Delta$ to the LP value. Therefore, there are at most $O(m/\Delta)$ -edges across all $S$. Since each -edge is visited at most once, listing all -edges takes at most $O(m/\Delta)$ time.
@@ 9.2 Rounding in Sublinear Model
On the other hand, for each vertex $v\in U_{r}$, the probability that it is included in the cluster of $r$ is at least $1-x_{rv}\geq\Delta$ for ++edges and $1-x_{rv}\cdot x_{rv}\geq\Delta$ for -edges. Thus, in expectation, we remove at least $\Delta|U_{r}|$ vertices from the current graph.